OpenStudy (anonymous):

A typical healthy person's blood pressure can be modeled by the periodic function below, where t is time (in seconds) and f(t) is in millimeters of mercury. Which one of 0.5, 0.8, or 1 appears to be the period of this function? f(t) = 26cos(2.5π t)+91

OpenStudy (anonymous):

I think 2.5 it's not frequency, it's angular velocity, as f(x)=Acos(wt) so it'd be 2*pi*f=w; f=w/2*pi ; f=2.5*pi/2*pi=1.25 and T=1/f so T=0.8 (s)

OpenStudy (anonymous):

I'm still lost and already got the problem wrong, but 0.8 was the correct answer.

OpenStudy (anonymous):

Got it wrong before putting this up here, I mean. I just wanted to know how to do it.

OpenStudy (anonymous):

so, what's the problem?

OpenStudy (anonymous):

tell me so I can helpl you

OpenStudy (anonymous):

Where did you get f in 2πf=w?

OpenStudy (anonymous):

Oh, wait, that's frequency, huh?

OpenStudy (anonymous):

f is frequency

OpenStudy (anonymous):

Ok, I get it. I just needed to read it about 3 times.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

sure? ok, yw. If you have more questions, just ask

OpenStudy (anonymous):

Will do!